# Renewal theory for Brownian motion with stochastically gated targets

### v1

@@ 5 Discussion
As previously shown for non-gated adsorption/desorption processes [45], the renewal approach is more complicated when considering interior traps rather then target boundaries, and this holds even in 1D. In order to illustrate this, consider a stochastically gated trap ${\mathcal{U}}=[a,b]$ in the interior of the half-line $\Omega=[0,\infty)$, see Fig. 7. (For an alternative renewal formulation of this particular FPT problem, which yields exact results for the FPT density, see Ref. [24].) The particle can now freely enter and exit $\mathcal{U}$} and is absorbed within $\mathcal{U}$} as soon as the gate opens. We would like to construct a renewal equation that relates the joint propagator $\rho_{k|k_{0}}$ to a marginal propagator $p$ for which the substrate is totally absorbing (non-gated). However, the latter would only represent sample paths that are restricted to the exterior of $\mathcal{U}$} since as soon as a particle reaches $\partial{\mathcal{U}}$ it is removed. Therefore, suppose that $\mathcal{U}$} is a partially adsorbing substrate with reactivity $\kappa_{0}$. If the gate is closed when the particle is adsorbed at a point $y\in{\mathcal{U}}$, then diffusion is immediately restarted from $y$ (without any shift in position). The solution of the original problem is recovered by taking the limit $\kappa_{0}\rightarrow\infty$, although keeping $\kappa_{0}$ finite is also of interest. In the latter case, the renewal equation takes the form
$\rho^{(\kappa_{0})}_{k|k_{0}}(x,t|x_{0})$ $=\Pi_{kk_{0}}(t)p^{(\kappa_{0})}(x,t|x_{0})$ (5.1)
$\quad+\int_{a}^{b}dy\int_{0}^{t}d\tau\,\rho^{(\kappa_{0})}_{k|-1}(x,t-\tau|y)\Pi_{-1,k_{0}}(\tau)p^{(\kappa_{0})}(y,\tau|x_{0}),$
with $p^{(\kappa_{0})}(x,t|x_{0})$ the propagator for a partially adsorbing substrate
$\frac{\partial p^{(\kappa_{0})}(x,t|x_{0})}{\partial t}$ $=D\frac{\partial^{2}p^{(\kappa_{0})}(x,t|x_{0})}{\partial x^{2}}-\kappa_{0}p^{(\kappa_{0})}(x,t|x_{0})\chi_{{\mathcal{U}}}(x),\quad 0<x<\infty,$ (5.2a)
$D\frac{\partial p^{(\kappa_{0})}(0,t|x_{0})}{\partial x}$ $=0,\quad p(x,0|x_{0})=\delta(x-x_{0}).$ (5.2b)
We have introduced the indicator function $\chi_{{\mathcal{U}}}(x)=1$ if $x\in{\mathcal{U}}$ and zero otherwise. In addition $J^{(\kappa_{0})}(y,t|x_{0})=\kappa_{0}p^{(\kappa_{0})}(y,t|x_{0})$ is the adsorption probability flux density at the point $y\in{\mathcal{U}}$. Comparison of equation (5.1) with the corresponding renewal equation (2.17) for a gated point-like boundary shows that the former involves an additional integration with respect to the spatial variable $y\in{\mathcal{U}}$, which cannot be eliminated by Laplace transforming with respect to $t$:
$\widetilde{\rho}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})=\widetilde{p}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})+\int_{a}^{b}dy\,\widetilde{\rho}_{k|-1}^{(\kappa_{0})}(x,s|y)\widetilde{p}^{(\kappa_{0})}_{-1|k_{0}}(y,s|x_{0}).$ (5.3)
Equation (5.3) is an example of a 1D Fredholm integral equation. One method for analysing such an equation is to perform a Neumann series expansion:
$\widetilde{\rho}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})$ $=\widetilde{p}^{(\kappa_{0})}_{k|k_{0}}(x,s|x_{0})+\int^{b}_{a}dy\,\widetilde{p}^{(\kappa_{0})}_{k|-1}(x,s|y)\widetilde{p}^{(\kappa_{0})}_{-1|k_{0}}(y,s|x_{0})$
$\hskip 28.45274pt+\int_{a}^{b}dy\int^{b}_{a}dy^{\prime}\widetilde{p}^{(\kappa_{0})}_{k|-1}(x,s|y)\widetilde{p}^{(\kappa_{0})}_{-1|-1}(y,s|y^{\prime})\widetilde{p}^{(\kappa_{0})}_{-1|k_{0}}(y^{\prime},s|x_{0})+\ldots,$ (5.4)
It can be shown that the series is uniformly convergent. However, a more practical issue is determining how many terms are required for a given level of numerical accuracy. (In higher spatial dimensions the Neumann series expansion could be combined with the spectral theory of D-to-N operators along analogous lines to Ref. [45].) It would also be interesting to understand the relationship between the series expansion of the corresponding 1D FPT density and the exact representation obtained in Ref. [24].
We end by noting one current limitation of our renewal theory for stochastic gating that does not occur for adsorption/desorption processes [45, 46]. In the latter case, the Laplace transformed renewal equations can still be solved explicitly when the waiting time density $\phi(\tau)$ for the time $\tau$ attached to a sticky wall is a non-exponential function. That is, one can construct non-Markovian models of desorption/absorption. On the other hand, the renewal equation (2.17), for example, involves the product $p_{k|k_{0}}(x,t|x_{0}):=\Pi_{kk_{0}}(t)p(x,t|x_{0})$, where $\Pi_{kk_{0}}(t)$ is the transition matrix for the switching gate. We relied on the fact that in the case of Markovian switching, $\Pi_{kk_{0}}(t)$ consists of a linear combination of exponential functions of time, so that we can express the Laplace transform $\widetilde{p}_{k|k_{0}}(x,s|x_{0})$ in terms of the corresponding marginal density $\widetilde{p}(x,s|x_{0})$. Incorporating non-Markovian models of a switching gate is more of a challenge.
@@
$\quad\sim\frac{1}{D}\int_{{\mathcal{C}}}[s\widetilde{p}(\mathbf{x},s|\mathbf{x}_{0})-\delta(\mathbf{x}-\mathbf{x}_{0})]d\mathbf{x},$ (A.2)
where $\Sigma_{\epsilon}$ denotes the flat end of the cylinder within $\Omega$. If $\mathbf{x}_{0}$ is in the bulk domain, then taking the limits $\epsilon,\sigma\rightarrow 0$ shows that the flux is continuous as it approaches the boundary, since the right-hand side of equation (A.2) vanishes. On the other hand, if $\mathbf{x}_{0}=\mathbf{z}\in\partial\Omega$ then taking the limits $\epsilon,\sigma\rightarrow 0$ gives
$D\nabla\widetilde{p}(\mathbf{y},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{y})-\lim_{\epsilon\rightarrow 0^{+}}D\nabla\widetilde{p}(\mathbf{y}-\epsilon\mathbf{n}(\mathbf{y}),s|\mathbf{z})\cdot\mathbf{n}(y)=-\overline{\delta}(\mathbf{y}-\mathbf{z}),$ (A.3)
@@
$J(\mathbf{y},s|\mathbf{z})\equiv-D\nabla\widetilde{p}(\mathbf{y},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{y})=\overline{\delta}(\mathbf{y}-\mathbf{z}),\quad\mathbf{y},\mathbf{z}\in\partial\Omega.$ (A.4)

### v2

@@ 5 Discussion
There is one limitation of the renewal theory for stochastically gated targets that does not occur for adsorption/desorption processes [45, 46]. In the latter case, the Laplace transformed renewal equations can still be solved explicitly when the waiting time density $\phi(\tau)$ for the time $\tau$ attached to a sticky wall is a non-exponential function. That is, one can construct non-Markovian models of desorption/absorption. On the other hand, the renewal equation (2.4gr), for example, involves the product $p_{k|k_{0}}(x,t|x_{0}):=\Pi_{kk_{0}}(t)p(x,t|x_{0})$, where $\Pi_{kk_{0}}(t)$ is the transition matrix for the switching gate. We relied on the fact that in the case of Markovian switching, $\Pi_{kk_{0}}(t)$ consists of a linear combination of exponential functions of time, so that we can express the Laplace transform $\widetilde{p}_{k|k_{0}}(x,s|x_{0})$ in terms of the corresponding marginal density $\widetilde{p}(x,s|x_{0})$. Incorporating non-Markovian models of a switching gate is more of a challenge.
@@
Another major challenge is extending the renewal theory to the case of stochastically-gated interior traps rather than target boundaries. Consider, for example, a stochastically gated trap ${\mathcal{U}}=[a,b]$ in the interior of the half-line $\Omega=[0,\infty)$, see Fig. 7. The particle can now freely enter and exit $\mathcal{U}$} and is absorbed within $\mathcal{U}$} as soon as the gate opens. We would like to construct a renewal equation that relates the joint propagator $\rho_{k|k_{0}}$ to a marginal propagator $p$ for which the substrate is totally absorbing (non-gated). However, the latter would only represent sample paths that are restricted to the exterior of $\mathcal{U}$} since as soon as a particle reaches $\partial{\mathcal{U}}$ it would be removed. Hence, this would exclude sample paths that start within the interior of the trap when the gate is closed. One way to proceed would be to treat $\mathcal{U}$} as a partially adsorbing substrate with reactivity $\kappa_{0}$ along analogous lines to Fig. 4(a) so that the particle only detects the state of the gate after it is adsorbed. Alternatively, one could treat the open substrate as partially absorbing along analogous lines to Fig. 4(b) and use the Green’s function method of Szabo et al. [6, 14]. However, both approaches lead to integral equations involving spatial integrals over the trapping region so that the analysis is considerably more involved, particularly with regards retaking the limit $\kappa_{0}\rightarrow\infty$. Recently Sher et al. [24] have developed a method that avoids the occurrence of spatial integrals. They derive a recursive relation between the FPT density for absorption to the corresponding FPT density for adsorption in the absence of gating by conditioning on whether the particle is located within or outside the trap when the gate reopens. This yields exact expressions for the Laplace transform of the FPT density for a 1D trap. However, extending the latter method to higher-dimensional traps is itself non-trivial.
[…]
Finally, note that another class of problems where the renewal formulation developed in this paper can be applied is diffusion across a stochastically gated interface within the interior of a domain. Consider the example of BM in $\mathbb{R}$} with an interface at $x=0$, see Fig. 8. Following along analogous lines to section 2, one can construct a renewal equation that sews together a sequence of BMs on the half-line with a totally adsorbing boundary at $x=0$. Now, however, each time the particle is adsorbed, the stochastic process is immediately restarted according to the following modified rule: if the gate is closed then diffusion restarts on the same side of the interface, whereas if the gate is open then diffusion restarts on either side of the interface with equal probability. Again, in order to ensure that diffusion restarts in a state that avoids immediate re-absorption, the particle is instantaneously shifted a distance $\epsilon$ from the origin whenever it reaches the interface. The renewal theory for stochastically gated interfaces is developed in a companion paper [55]. For a corresponding theory of semi-permeable interfaces based on so-called snapping out BM see Ref. [54].
@@
$\quad\sim\frac{1}{D}\int_{{\mathcal{C}}}[s\widetilde{p}(\mathbf{x},s|\mathbf{x}_{0})-\delta(\mathbf{x}-\mathbf{x}_{0})]d\mathbf{x},$ (A.1b)
where $\Sigma_{\epsilon}$ denotes the flat end of the cylinder within $\Omega$. If $\mathbf{x}_{0}$ is in the bulk domain, then taking the limits $\epsilon,\sigma\rightarrow 0$ shows that the flux is continuous as it approaches the boundary, since the right-hand side of equation (A.1b) vanishes. On the other hand, if $\mathbf{x}_{0}=\mathbf{z}\in\partial\Omega$ then taking the limits $\epsilon,\sigma\rightarrow 0$ gives
$D\nabla\widetilde{p}(\mathbf{y},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{y})-\lim_{\epsilon\rightarrow 0^{+}}D\nabla\widetilde{p}(\mathbf{y}-\epsilon\mathbf{n}(\mathbf{y}),s|\mathbf{z})\cdot\mathbf{n}(y)=-\overline{\delta}(\mathbf{y}-\mathbf{z}),$ (A.1c)
@@
$J(\mathbf{y},s|\mathbf{z})\equiv-D\nabla\widetilde{p}(\mathbf{y},s|\mathbf{z})\cdot\mathbf{n}(\mathbf{y})=\overline{\delta}(\mathbf{y}-\mathbf{z}),\quad\mathbf{y},\mathbf{z}\in\partial\Omega.$ (A.1d)